40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 26 17:09:15 2012, Den, Update, SUS, TT 
    Reply  Thu Sep 27 15:21:04 2012, Steve, Update, SUS, TT height adaptor & PTICH damping 09271201.PDFIMG_1656.JPGIMG_1654.JPG
       Reply  Fri Oct 12 10:39:45 2012, Steve, Update, SUS, TT PITCH damping IMG_1709.JPG
          Reply  Fri Oct 12 14:01:33 2012, Koji, Update, SUS, TT PITCH damping 
             Reply  Fri Oct 12 16:59:26 2012, steve, Update, SUS, TT PITCH damping 
       Reply  Wed Oct 17 08:30:49 2012, Steve, Update, SUS, PTICH damping needed 
          Reply  Wed Oct 17 11:15:05 2012, Den, Update, SUS, PITCH damping needed pitch_damping.png
             Reply  Thu Oct 18 02:42:38 2012, Den, Update, SUS, PITCH damping needed 7x
                Reply  Fri Oct 19 18:30:07 2012, Den, Update, SUS, PITCH damping needed damp.pngpitch.png
                   Reply  Sat Oct 20 01:23:56 2012, Den, Update, SUS, PITCH damping needed P1010059.jpgP1010058.jpgP1010060.jpg
                      Reply  Sat Oct 20 20:37:55 2012, Den, Update, SUS, PITCH damping needed DSC_4778.JPGpitch.JPGyaw.JPGDSC_4776.JPG
                      Reply  Mon Oct 22 11:26:12 2012, rana, Update, SUS, PITCH damping needed 
                         Reply  Tue Oct 23 01:46:53 2012, Den, Update, SUS, PITCH damping needed Mirror_Holder_ECD_NEW_DRAWING.PDFTT_DAMPING_DRAWING.PDF
                            Reply  Tue Oct 23 17:12:30 2012, Den, Update, SUS, PITCH damping needed simple_drawing.PDF
                               Reply  Wed Oct 24 01:02:10 2012, Koji, Update, SUS, PITCH damping needed 
                                  Reply  Mon Oct 29 18:14:55 2012, Den, Update, SUS, PITCH damping needed DSC_4787.JPGDSC_4791.JPG
                            Reply  Tue Oct 23 18:18:29 2012, Jenne, Update, SUS, PITCH damping needed 
Message ID: 7640     Entry time: Mon Oct 29 18:14:55 2012     In reply to: 7604
Author: Den 
Type: Update 
Category: SUS 
Subject: PITCH damping needed 

 

 We've received all parts that we need for eddy current damping. I've made an estimate of Q with dirty tip-tilt. It looks fine (Q~1)

We need to check ring magnets for vacuum compatibility. Bob start baking on Friday.

DSC_4787.JPG     DSC_4791.JPG

ELOG V3.1.3-